14.60g sample of an unknown compound, composed only of carbon, hydrogen, and oxygen, produced 28.6g of CO2 and 14.6g of H2O in a combustion analysis. What is the mass percent composition of oxygen of the unknown compound?
1
Expert's answer
2019-10-02T08:03:08-0400

Mass of carbon +mass of hydrogen+mass of oxygen = 14.6"\ \ \ \ \ \ ........(1)"

"Mass\ of\ C\ in\ CO_2=\frac{12}{44}\times 28.6=7.8\ g"

"Mass\ of\ H\ in \ H_2O=\frac{2}{18}\times14.6=1.6222\ g"

these carbon and hydrogen came only from the unknown sample

So,

Mass of oxygen in unknown sample = "14.6-7.8-1.6222=5.1778\ g"

Percentage composition of oxygen in sample = "\frac{5.1778}{14.6}\times 100=35.464\%"